Assume that the random variable X is normally distributed, with mean μ = 70 and standard deviation σ = 13.

Find P(X ≤ 65) =
Answer #1

Solution :

Given that ,

mean\mu =  = 70

standard deviation = \sigma =13

P(x  \leq65)

= P[(x - \mu ) / \sigma \leq (65-70) / 13]

= P(z \leq -0.38 )

Using z table,

=0.3520